OpenStudy (anonymous):

The variable t represents the temperature at 6 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Washington dropped 19°F between 6 p.m. and 10 p.m. The temperature at 10 p.m. was –12°F. What was the temperature at 6 p.m.?

OpenStudy (anonymous):

A. 12-19=t B. 19+12=t C. t-19=-12 D. 19-t=-12